a cone has diameter of 10 and height of 5
Answer by avani(25)

The volume of a cone is 1/3πr^2h
where r is the radius,h is the height, the radius is d/2=10/2 = 5
Thus,

The volume of a cone is 1/3πr^2h = 1/3 * π * 5^2 * 5 = 130.90 units